Oil painting frame- The perimeter of the ornate frame of an oil painting is 146in. The width is 5in. Less than the length. Find the demotions
Answer by Finavon(81)

l & w for length and width
2l+2w=146
l=w+5
substitute for l
2w+10+2w=146
4w=136
w=34
l=39
Frame is 34 in by 39 in
Check: perimeter = 2*(34+39) = 2*73 = 146